# A Novel Deep Hybrid Framework with Ensemble-Based Feature Optimization for Robust Real-Time Human Activity Recognition

### v2

@@ 5.2 Evaluation Metrics
To comprehensively evaluate the performance of the proposed framework, several quantitative metrics were employed. Classification accuracy (%) was computed on the test set to assess the overall predictive capability of the model by Equation 32. In addition, the precision, recall, F1-score, were used to provide a more balanced evaluation, particularly in the presence of class imbalance by Equations 3335. The number of selected features was recorded to reflect the dimensionality reduction achieved through the FS process. To assess computational efficiency, the average inference time (in milliseconds) was measured over five independent runs. Furthermore, memory usage (in KB) of the selected features was analyzed to evaluate the memory footprint of the model. These metrics together offer a holistic view of the classification performance, computational efficiency, and practicality of the proposed framework.
@@ 5.3 Performance Summary
Table 5 presents a comparative analysis of the proposed ADFSA FS technique against conventional techniques including PCA and RFE across four widely used classifiers: SVM, RF, KNN, and LogReg. The evaluation is based on three key metrics: cross-validation accuracy, number of selected features, and average inference time over five runs. The ADFSA method consistently achieves the highest accuracy across all classifiers, reaching up to 99.65% for RF, while significantly reducing the feature set to only seven features. This demonstrates the method’s capability to retain highly discriminative features while discarding redundant or irrelevant ones. Additionally, ADFSA achieves the lowest inference times in all cases, indicating its suitability for real-time human activity recognition applications. The combined improvements in accuracy, feature compactness, and computational efficiency clearly highlight the superiority of the proposed ensemble fitness function over traditional FS techniques.
@@
Figures 10 and 11 illustrate a comprehensive comparative evaluation of different FS methods applied to SVM and RF classifiers, respectively. The results clearly highlight the superior performance of the proposed ADFSA technique in all metrics considered. For the SVM classifier, ADFSA achieves the highest accuracy of 99.58%, outperforming traditional methods such as PCA and RFE, which yield 97.79% and 97.73%, respectively. Notably, ADFSA also uses only 7 features, significantly fewer than PCA (50) and the baseline (128), demonstrating its efficiency in dimensionality reduction. Additionally, it reports the lowest inference time of 0.196 ms, indicating suitability for real-time scenarios.
@@
Figures 12 and 13 further validate the effectiveness of the proposed ADFSA FS technique when applied to KNN and LogReg classifiers. For the KNN classifier (Figure 12), ADFSA achieves the highest accuracy of 99.58%, outperforming PCA (97.54%) and RFE (97.44%). More importantly, it reduces the number of features to just 7, compared to 50 for PCA/RFE and 128 in the baseline. Inference time also shows a significant improvement, dropping to 6.958 ms, whereas PCA and RFE report 8.033 ms and 17.676 ms, respectively. This performance reinforces the scalability of ADFSA even in non-parametric models like KNN.
@@ 5 Soccer Juggling
5.4 Analysis of Memory Usage and Training Time
@@ 5.5 Other Performance Matrices
Furthermore, Figures 20-23 present a comparative analysis of the performance metrics, including F1-score, Precision, and Recall, for four classifiers: SVM, RF, KNN, and LogReg, respectively. Each figure includes grouped bar charts that visualize the impact of four FS methods: No FS, PCA, RFE, and the proposed ADFSA technique.
@@ 5.6 Ablation Study
Figure 25 and 26 illustrates the impact of different fitness function configurations on the optimizer’s weight update behavior and FS performance. In the first setting, Fig. 25(a), where neither uniqueness nor complexity penalties are applied, the optimizer enters a stagnation phase after just 6 generations, resulting in the selection of 52 features. This early stagnation indicates poor exploration capability and a tendency to converge prematurely. In the second configuration, Fig. 25(b), which includes only the complexity penalty, the optimizer shows a modest improvement by stagnating after 12 generations and selecting 64 features, suggesting some control over model complexity but still lacking diversity in the solution space. In contrast, the third configuration, Fig. 26 uses the complete fitness function with both uniqueness and complexity components. This setup allows the optimizer to utilize the full range of generations without premature stagnation and leads to a significantly reduced and optimal feature subset of only 7 features. These results demonstrate the importance of balancing exploration and exploitation through well-designed fitness objectives in FS tasks.

### v3

@@ 5.3 Evaluation Metrics
To comprehensively evaluate the performance of the proposed framework, several quantitative metrics were employed. Classification accuracy (%) was computed on the test set to assess the overall predictive capability of the model by Equation 32. In addition, the precision, recall, F1-score, were used to provide a more balanced evaluation, particularly in the presence of class imbalance by Equations 3335. The number of selected features was recorded to reflect the dimensionality reduction achieved through the FS process. To assess computational efficiency, the average inference time (in milliseconds) was measured over five independent runs. Furthermore, the memory usage (in KB) of the selected features was analyzed to evaluate the memory footprint of the model. These metrics together offer a holistic view of the classification performance, computational efficiency, and practicality of the proposed framework.
@@ 5.4 Performance Summary
Table 5 presents a comparative analysis of the proposed ADFSA FS technique against conventional techniques including PCA and RFE across four widely used classifiers: SVM, RF, KNN, and LogReg. The evaluation is based on three key metrics: cross-validation accuracy, number of selected features, and average inference time over five runs. The ADFSA method consistently achieves the highest accuracy across all classifiers, reaching up to 99.65% for RF, while significantly reducing the feature set to only seven features. This demonstrates the method’s capability to retain highly discriminative features while discarding redundant or irrelevant ones. Additionally, ADFSA achieves the lowest inference times in all cases, indicating its suitability for real-time human activity recognition applications. The combined improvements in accuracy, feature compactness, and computational efficiency clearly highlight the superiority of the proposed dynamic composite fitness function over traditional FS techniques.
@@
Figures 8 and 9 illustrate a comprehensive comparative evaluation of different FS methods applied to SVM and RF classifiers, respectively. The results clearly highlight the superior performance of the proposed ADFSA technique in all metrics considered. For the SVM classifier, ADFSA achieves the highest accuracy of 99.58%, outperforming traditional methods such as PCA and RFE, which yield 97.79% and 97.73%, respectively. Notably, ADFSA also uses only 7 features, significantly fewer than PCA (50) and the baseline (128), demonstrating its efficiency in dimensionality reduction. Additionally, it reports the lowest inference time of 0.196 ms, indicating suitability for real-time scenarios.
@@
Figures 10 and 11 further validate the effectiveness of the proposed ADFSA FS technique when applied to KNN and LogReg classifiers. For the KNN classifier (Figure 10), ADFSA achieves the highest accuracy of 99.58%, outperforming PCA (97.54%) and RFE (97.44%). More importantly, it reduces the number of features to just 7, compared to 50 for PCA/RFE and 128 in the baseline. Inference time also shows a significant improvement, dropping to 6.958 ms, whereas PCA and RFE report 8.033 ms and 17.676 ms, respectively. This performance reinforces the scalability of ADFSA even in non-parametric models like KNN.
@@ 5 Soccer Juggling
5.5 Memory Usage and Training Time Analysis
@@ 5.6 Other Performance Matrices
Furthermore, Figures 18-21 present a comparative analysis of the performance metrics, including F1-score, Precision, and Recall, for four classifiers: SVM, RF, KNN, and LogReg, respectively. Each figure includes grouped bar charts that visualize the impact of four FS methods: No FS, PCA, RFE, and the proposed ADFSA technique.
Figure 18: Performance of SVM using different FS Methods.
@@ 5.7 Ablation Study
Figure 23 and 24 illustrates the impact of different fitness function configurations on the optimizer’s weight update behavior and FS performance. In the first setting, Fig. 23(a), where neither uniqueness nor complexity penalties are applied, the optimizer enters a stagnation phase after just 6 generations, resulting in the selection of 52 features. This early stagnation indicates poor exploration capability and a tendency to converge prematurely. In the second configuration, Fig. 23(b), which includes only the complexity penalty, the optimizer shows a modest improvement by stagnating after 12 generations and selecting 64 features, suggesting some control over model complexity but still lacking diversity in the solution space. In contrast, the third configuration, Fig. 24 uses the complete fitness function with both uniqueness and complexity components. This setup allows the optimizer to utilize the full range of generations without premature stagnation and leads to a significantly reduced and optimal feature subset of only 7 features. These results demonstrate the importance of balancing exploration and exploitation through well-designed fitness objectives in FS tasks.